Could you help me solve this question: Find the probability of randomly selecting a point inside the region bounded by segments joining the points (1,1), (-1,1), (-1,-1), and (1,-1) such that the sum of the coordinates is less than one-half. So far I have plotted the points but I don't know what to do! Please help, my grade is really struggling and this question will be on my test. Thanks!
Answer by sudhanshu_kmr(1152)


Area bounded by points (1,1), (-1,1), (-1,-1), and (1,-1) = 4 unit sq.

Area bounded by points that will satisfy the condition i.e (1/4,1/4), (-1/4,1/4),
(-1/4,-1/4), and (1/4,-1/4) = 1/4 unit sq.


required probability = 1/16

if any doubt, you are welcome to contact...